OpenStudy (anonymous):

Without graphing, is the system independent, dependent, or inconsistent? A. Independent B. Depended C. Inconsistent

OpenStudy (anonymous):

Please Help! An explanation too

OpenStudy (anonymous):

from equation 1st put value of y in 2nd to get equation in x

OpenStudy (anonymous):

What is the value of y?

OpenStudy (anonymous):

Are you telling me to combine both equations?

OpenStudy (anonymous):

both the equation are same

OpenStudy (anonymous):

So then what do I do?

OpenStudy (anonymous):

A consistent independent system always has a single solution.

OpenStudy (anonymous):

since both eqution are same so they overllap and infinite many solution .

OpenStudy (nurali):

This explains whether dependent, independent or inconsistent http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

OpenStudy (anonymous):

Does this mean they are inconsistent ?

OpenStudy (anonymous):

not if there is no solution then incosistent .

OpenStudy (anonymous):

there are infinitely many solutions and the system is called dependent.

OpenStudy (anonymous):

Ok, thanks for the help! But now how would I know if something is independent?

OpenStudy (anonymous):

mark out that one solution = independent or consistent no solution = in cosistent infinitaly many solution = dependent

OpenStudy (anonymous):

Really helpful! Thanks Can I post another?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

What would be the equivalent system to help me solve it?

OpenStudy (anonymous):

i dont get but solve them using value of y from first equation put in 2nd one .

OpenStudy (anonymous):

So basically, 3x - y + -2x?

OpenStudy (anonymous):

I combined both equations but I don't think it's right though Can you explain?

OpenStudy (anonymous):

not from 1st y=-2x now put in 2nd 3x-(-2x)=10

OpenStudy (anonymous):

Now when you say 'solve', I need to subtract something from both sides?

OpenStudy (anonymous):

not 3x+2x=10 5x=10 x=2

OpenStudy (anonymous):

Oh so we combined 'like terms' then

OpenStudy (anonymous):

If x=2, then the answer must be C. (2, -4) since that is the only option with a positive two in the x side, right?

OpenStudy (anonymous):

to solve mean from one equation put value of either of y or x in term of otherone now other equation has only x now solve for x and get answer..

OpenStudy (anonymous):

So C isn't the answer?

OpenStudy (anonymous):

no c is ans

OpenStudy (anonymous):

since on solving them we get x=2 and y=-4

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

from 1st y=-2x-1 put in second equation

OpenStudy (anonymous):

y = 2x - 1- 4x ? or y = 2x-1+4x ?

OpenStudy (anonymous):

-4x-2(-2x-1)=-1

OpenStudy (anonymous):

Oh, sorry, I completely forgot about the parenthesis

OpenStudy (anonymous):

-6x -3 = 1?

OpenStudy (anonymous):

I got that answer by combining 'like terms'

OpenStudy (anonymous):

not here -4x+4x+2=1 2=1

OpenStudy (anonymous):

now u think what is this this the case where there is no solution since line are parallal